# Maximum Preference Optimization with Importance Sampling

### 27 Dec 2023 v1

@@ 4 Method
In this work, we use the same objective of IPO. Unlike IPO’s loss, we derive a off-policy algorithm for preference maximization from a importance sampling policy optimization view, and add a intuitive sample loss to make KL-regularation term work.
4.1 Total Preference Optimization using Importance Sampling
First we define $x,y_{w},y_{l}$ as state, and $\mathcal{A}_{x,y_{w},y_{l}}=\{y_{w}\succ y_{l},y_{l}\succ y_{w}\}$ as action set, action as $\mathfrak{a}\in\mathcal{A}$ and define the reward of action $\mathfrak{a}$ as $r^{p}(\mathfrak{a})$, so
@@ 4.1 Total Preference Optimization using Importance Sampling
$\{\langle x,y_{w},y_{l},I\rangle\}\rightarrow\{\langle\underbrace{(x,y_{w},y_{l})}_{\textnormal{state}},\underbrace{(y_{w}\succ y_{l})}_{\textnormal{action}},\underbrace{(I)}_{\textnormal{reward}}\rangle,\langle\underbrace{(x,y_{w},y_{l})}_{\textnormal{state}},\underbrace{(y_{l}\succ y_{w})}_{\textnormal{action}},\underbrace{(1-I)}_{\textnormal{reward}}\rangle\},$ (21)
and define the new dataset as $\bar{\mathcal{D}}^{p}$. Because both actions appear at the same time from $\bar{\mathcal{D}}^{p}$, we can define the policy generating $\bar{\mathcal{D}}^{p}$ as:
$\bar{\pi}^{p}(\mathfrak{a}|x,y_{w},y_{l})=\bar{\pi}^{p}(\mathfrak{a}|x)=1/2,\forall\mathfrak{a}\in\mathcal{A}.$ (22)
Now we can formulate the distribution to generate $\bar{\mathcal{D}}^{p}$ as:
@@
and its easily to get $R(\bar{\pi}^{p})=1/2$.
Express the preferences maximization objective in reinforcement learning format:
$\max_{\pi^{p}_{\theta}}R(\pi^{p}_{\theta})$ (25)
@@
$\underset{\begin{subarray}{c}x\sim\rho\\ \langle y_{w},y_{l}\rangle\sim\mu(\cdot|x)\\ \mathfrak{a}\sim\pi^{p}_{\theta}(\cdot|x)\end{subarray}}{\mathbb{E}}[r^{p}(\mathfrak{a}|x)\log\pi^{p}_{\theta}(\mathfrak{a}|x)]$ (26)
Typically this gradient needs to be estimated from samples continually collected by $\pi^{p}_{\theta}$, which is data-inefficient. Using importance sampling, gradient 26 be expressed as:
$\underset{\begin{subarray}{c}x\sim\rho\\ \langle y_{w},y_{l}\rangle\sim\mu(\cdot|x)\\ \mathfrak{a}\sim\bar{\pi}^{p}(\cdot|x)\end{subarray}}{\mathbb{E}}[\frac{\pi^{p}_{\theta}(\mathfrak{a}|x)}{\bar{\pi}^{p}(\mathfrak{a}|x)}r^{p}(\mathfrak{a}|x)\log\pi^{p}_{\theta}(\mathfrak{a}|x)]$ (27)
Gradient 27 can be calculated offline, and the corresponding algorithm is off-policy. According to equation 22, $\bar{\pi}^{p}(\mathfrak{a}|x)=1/2,\forall\mathfrak{a}\in\mathcal{A}$, so gradient 26 and 27 are identical.
Theorem 1
Objective 25 (which is identical to preferences maximization term of IPO) can be directly optimized using off-policy gradient ascent algorithm.
@@ 4.2 Off-policy Preference Learning under KL-regularation
Define the corresponding point-wise policy for $\pi^{p}_{\theta}$ as $\pi^{s}_{\theta}$, we use Bradley-Terry model to represent $\pi^{p}_{\theta}$ as:
@@
KL-regularized preferences maximization objective 25 can be expressed as:
$\begin{gathered}\max_{\pi^{p}_{\theta},\pi^{s}_{\theta}}R(\pi^{p}_{\theta})-\tau\mathbb{D}_{\textnormal{KL}}[\pi^{s}_{\theta}||\pi_{\textnormal{ref}}],\end{gathered}$ (29)
which includes two expectations over different distributions:
[…]
From theorem 1, preference maximization can be directly be solved with off-policy gradient descent. But the KL-regularization term needs to collect samples from $\pi^{s}_{\theta}(\cdot|x)$, which is on-policy. Additionally, to optimize KL-regularization with stochastic samples, it may ignore the regularization on $\pi_{\textnormal{ref}}(y_{w}|x),\pi_{\textnormal{ref}}(y_{l}|x)$.
Off-policy KL-regularation optimization
Minimize $\mathbb{D}_{\textnormal{KL}}[\pi^{s}_{\theta}||\pi_{\textnormal{ref}}]$ needs on-policy samples collection, which is data-inefficient. To solve the problem, we use use $\mathbb{D}_{\textnormal{KL}}[\pi_{\textnormal{ref}}||\pi^{s}_{\theta}]$ as the KL-regularization term. Because $\pi_{\textnormal{ref}}$ is fixed, we can generate samples $\mathcal{D}_{\pi_{\textnormal{ref}}(\cdot|x)}$ conditioned on $x\sim\rho$ in advance. Then when optimize $\mathbb{D}_{\textnormal{KL}}[\pi_{\textnormal{ref}}||\pi^{s}_{\theta}]$, we sample datas from $\mathcal{D}_{\pi_{\textnormal{ref}}(\cdot|x)}$.
Explicit regulariztion on $\pi_{\textnormal{ref}}(y_{w}|x),\pi_{\textnormal{ref}}(y_{l}|x)$
Optimize KL-regularization may using stochastic samples may ignore regularization on $\pi_{\theta}(y_{w}|x),\pi_{\theta}(y_{l}|x)$. To solve this problem, we add a KL-regularation term for preference probability.
Define $\pi^{p}_{\textnormal{ref}}(y_{w}\succ y_{l}|x)=\sigma(\log\pi_{\textnormal{ref}}(y_{w}|x)-\log\pi_{\textnormal{ref}}(y_{l}|x))$ , then
$\begin{split}\mathbb{D}_{\textnormal{KL}}[\pi^{p}_{\theta}||\pi^{p}_{\textnormal{ref}}]=&\underset{\begin{subarray}{c}x\sim\rho\\ y_{w},y_{l}\sim\mu(\cdot|x)\\ \mathfrak{a}\sim\pi^{p}_{\theta}(\cdot|x)\end{subarray}}{\mathbb{E}}[\log\frac{\pi^{p}_{\theta}(\mathfrak{a}|x)}{\pi^{p}_{\textnormal{ref}}(\mathfrak{a}|x)}]\\ =&\underset{\begin{subarray}{c}x\sim\rho\\ y_{w},y_{l}\sim\mu(\cdot|x)\end{subarray}}{\mathbb{E}}[\sum_{\mathfrak{a}\in\mathcal{A}}\pi^{p}_{\theta}(\mathfrak{a}|x)\log\frac{\pi^{p}_{\theta}(\mathfrak{a}|x)}{\pi^{p}_{\textnormal{ref}}(\mathfrak{a}|x)}].\end{split}$ (31)
@@
Combine preference optimization term, point-wise policy KL-regularation term and pair-wise preference KL-regularation term, we get the final objective:
$\max_{\pi^{p}_{\theta},\pi^{s}_{\theta}}=\underset{x\sim\rho}{\mathbb{E}}\Big{[}\underset{y_{w},y_{l}\sim\mu(\cdot|x)}{\mathbb{E}}\big{\{}R(\pi^{p}_{\theta}(\cdot|x,y_{w},y_{l}))-\beta\mathbb{D}_{\textnormal{KL}}[\pi^{p}_{\theta}||\pi^{p}_{\textnormal{ref}}]\big{\}}-\tau\mathbb{D}_{\textnormal{KL}}[\pi_{\textnormal{ref}}||\pi^{s}_{\theta}]\Big{]}.$ (32)
With this objective, we define MPO loss on dataset $\bar{\mathcal{D}}^{p}$ and $\mathcal{D}_{\pi_{\textnormal{ref}}}$:
$\mathcal{L}_{\textnormal{MPO}}=\underset{\begin{subarray}{c}<x,y_{w},y_{l},\mathfrak{a},I>\sim\bar{\mathcal{D}}^{p}\\ \end{subarray}}{\mathbb{E}}\Big{[}-I\pi^{p}_{\theta}(\mathfrak{a}|x)+\beta\pi^{p}_{\theta}(\mathfrak{a}|x)\log\frac{\pi^{p}_{\theta}(\mathfrak{a}|x)}{\pi^{p}_{\textnormal{ref}}(\mathfrak{a}|x)}-\tau\underset{y\sim\mathcal{D}_{\textnormal{ref}}(\cdot|x)}{\mathbb{E}}[\log{\pi^{s}_{\theta}(y|x)}]\Big{]}.$ (33)
This loss is very simple and intuitive, term $-I\pi^{p}_{\theta}(\mathfrak{a}|x)$ try to maximize preferences, $\beta$ controls the strength of the KL regularization between pair-wise preference generating distribution, and $\tau$ controls the strength of the KL regularization between the point-wise generating distribution.
Eliminate $\pi_{\textnormal{ref}}$ after sample generation
When generating $\mathcal{D}_{\textnormal{ref}}(\cdot|x)$ and calculating $\log\pi^{p}_{\textnormal{ref}}(\mathfrak{a}|x)$ in advance, we can optimize $\mathcal{L}_{\textnormal{MPO}}$ without $\pi_{\textnormal{ref}}$. Which means we can optimize $\mathcal{L}_{\textnormal{MPO}}$ from $\pi_{\textnormal{ref}}$, without any modification of the model structure but loss function.

### 31 Dec 2023 v2

@@ 4 Method
In this work, we combine the preference maximization term of IPO’s loss and modified regularization term of RLHF’s loss. Unlike IPO, which derive the off-policy from preference maximization under KL-regularization, we formulate preference maximization as a reward maximization problem in reinforcement learning setting, and derive an off-policy objective from an importance sampling based policy optimization view, but without the help of KL-regularation. Then we combine the off-policy reward maximization objective with modified regularization terms of RLHF’s ’PPO-ptx’ objective, which makes the KL-regularization truly effective. We call the algorithm Maximum Preference Optimization (MPO). The final objective of MPO bears resemblance to RLHF’s objective, and likes IPO, MPO is off-policy.
4.1 Preference(reward) Maximization with Importance Sampling
We define preference as reward, and formalize preference maximization as reward maximization problem in reinforcement learning setting. Define $x,y_{w},y_{l}$ as state, and $\mathcal{A}_{x,y_{w},y_{l}}=\{y_{w}\succ y_{l},y_{l}\succ y_{w}\}$ as action set, action as $\mathfrak{a}\in\mathcal{A}$ and define the reward of action $\mathfrak{a}$ as $r^{p}(\mathfrak{a}|x,y_{w},y_{l})$, which is the preference probability. We simplify $r^{p}(\mathfrak{a}|x,y_{w},y_{l})$ as $r^{p}(\mathfrak{a}|x)$, so:
@@ 4.1 Preference(reward) Maximization with Importance Sampling
$\{\langle x,y_{w},y_{l},I\rangle\}\rightarrow\{\langle\underbrace{(x,y_{w},y_{l})}_{\text{state}},\underbrace{(y_{w}\succ y_{l})}_{\text{action}},\underbrace{(I)}_{\text{reward}}\rangle,\langle\underbrace{(x,y_{w},y_{l})}_{\text{state}},\underbrace{(y_{l}\succ y_{w})}_{\text{action}},\underbrace{(1-I)}_{\text{reward}}\rangle\},$ (21)
and define the converted dataset as $\bar{\mathcal{D}}^{p}$. Because both actions appear at the same time from $\bar{\mathcal{D}}^{p}$, we can define the policy generating $\bar{\mathcal{D}}^{p}$ as:
$\bar{\pi}^{p}(\mathfrak{a}|x,y_{w},y_{l})=1/2,\forall\mathfrak{a}\in\mathcal{A}.$ (22)
Given the state $\langle x,y_{w},y_{l}\rangle$, we simplify $\bar{\pi}^{p}(\mathfrak{a}|x,y_{w},y_{l})$ as $\bar{\pi}^{p}(\mathfrak{a}|x)$. Now we can formulate the distribution to generate $\bar{\mathcal{D}}^{p}$ as:
@@
and it’s easy to get $R(\bar{\pi}^{p})=1/2$.
Express the preference maximization objective in reinforcement learning context:
$\max_{\pi^{p}_{\theta}}R(\pi^{p}_{\theta}).$ (25)
Typically, the gradient of the objective 25 needs to be estimated from samples continually collected by $\pi^{p}_{\theta}$, which is data-inefficient. But, for preference maximization, we can directly estimate gradient from dataset $\bar{\mathcal{D}}^{p}$, which is off-policy.
Theorem 1.
Preference(Reward) maximization objective 25 (which is identical to preferences maximization term of IPO) can be directly optimized using off-policy gradient ascent algorithm.
Proof.
@@
$\nabla R(\pi^{p}_{\theta})=\underset{\begin{subarray}{c}x\sim\rho\\ \langle y_{w},y_{l}\rangle\sim\mu(\cdot|x)\\ \mathfrak{a}\sim\pi^{p}_{\theta}(\cdot|x)\end{subarray}}{\mathbb{E}}[r^{p}(\mathfrak{a}|x)\log\pi^{p}_{\theta}(\mathfrak{a}|x)].$ (26)
Using importance sampling, gradient 26 be expressed as:
$\nabla R(\pi^{p}_{\theta})=\underset{\begin{subarray}{c}x\sim\rho\\ \langle y_{w},y_{l}\rangle\sim\mu(\cdot|x)\\ \mathfrak{a}\sim\bar{\pi}^{p}(\cdot|x)\end{subarray}}{\mathbb{E}}[\frac{\pi^{p}_{\theta}(\mathfrak{a}|x)}{\bar{\pi}^{p}(\mathfrak{a}|x)}r^{p}(\mathfrak{a}|x)\log\pi^{p}_{\theta}(\mathfrak{a}|x)]$ (27)
Gradient 27 can be calculated offline, and the corresponding algorithm is off-policy. According to equation 22, $\bar{\pi}^{p}(\mathfrak{a}|x)=1/2,\forall\mathfrak{a}\in\mathcal{A}$, so gradient 26 and 27 are identical. ∎
@@ 4.2 Off-policy Preference Learning under KL-regularation
Define the corresponding point-wise policy for $\pi^{p}_{\theta}$ as $\pi^{s}_{\theta}$, we use Bradley-Terry model to approximate $\pi^{p}_{\theta}$ as:
@@
Then, reward maximization objective 25 can be expressed:
$\max_{\pi^{s}_{\theta}}R(\pi^{p}_{\theta}),$ (29)
which means we optimize $\pi^{s}_{\theta}$ to maximize the preference reward for corresponding policy $\pi^{p}_{\theta}$.
Let $\pi^{s}_{\text{ref}}=\pi_{\text{ref}}$, like RLHF’s ’PPO-ptx’ objective 11, KL-regularized preferences maximization objective 25 can be expressed as:
$\begin{gathered}\max_{\pi^{s}_{\theta}}R(\pi^{p}_{\theta})-\beta\mathbb{D}_{\text{KL}}[\pi^{s}_{\theta}||\pi^{s}_{\text{ref}}]+\gamma\underset{x\sim\mathcal{D}_{\text{pretrain}}}{\mathbb{E}}[\log\pi^{s}_{\theta}(x)]\end{gathered}.$ (30)
From theorem 1, preference maximization term $R(\pi^{p}_{\theta})$ can be directly be solved with off-policy policy gradient method. Pretraining data regularization term $\mathbb{E}_{s\sim\mathcal{D}_{\text{pretrain}}}\log\pi^{s}_{\theta}(x)$ can also be solved with offline data. But the KL-regularization term $\mathbb{D}_{\text{KL}}[\pi^{s}_{\theta}||\pi^{s}_{\text{ref}}]$ needs to collect samples from $\pi^{s}_{\theta}(\cdot|x)$, which is on-policy.
Off-policy KL-regularation on reference policy $\pi^{s}_{\text{ref}}$
Minimize $\mathbb{D}_{\text{KL}}[\pi^{s}_{\theta}||\pi^{s}_{\text{ref}}]$ needs on-policy samples collection, which is data-inefficient. To solve the problem, we replace $\mathbb{D}_{\text{KL}}[\pi^{s}_{\theta}||\pi^{s}_{\text{ref}}]$ with $-\mathbb{E}_{\langle x,y\rangle\sim\mathcal{D}_{\text{ref}}}[\log\pi^{s}_{\theta}(y|x)]$. Like pretraining data regularization, replaced regularization can be computed with offline data.
@@
With the modified regularization on $\pi^{s}_{\text{ref}}$, we get the final objective of MPO:
$\max_{\pi^{s}_{\theta}}R(\pi^{p}_{\theta})-\beta\underset{\langle x,y\rangle\sim\mathcal{D}_{\text{ref}}}{\mathbb{E}}[\log\pi^{s}_{\theta}(y|x)]-\gamma\underset{x\sim\mathcal{D}_{\text{pretrain}}}{\mathbb{E}}[\log\pi^{s}_{\theta}(x)]$ (31)
With this objective, we define empirical MPO loss on dataset $\bar{\mathcal{D}}^{p}$,$\mathcal{D}_{\text{ref}}$ and $\mathcal{D}_{\text{pretrain}}$:
$\mathcal{L}_{\text{MPO}}=\underset{\begin{subarray}{c}<x,y_{w},y_{l},\mathfrak{a},I>\sim\bar{\mathcal{D}}^{p}\\ \end{subarray}}{\mathbb{E}}[-I\pi^{p}_{\theta}(\mathfrak{a}|x)]+\beta\underset{\langle x,y\rangle\sim\mathcal{D}_{\text{ref}}}{\mathbb{E}}[\log\pi^{s}_{\theta}(y|x)]+\gamma\underset{y\sim\mathcal{D}_{\text{pretrain}}}{\mathbb{E}}[\log{\pi^{s}_{\theta}(y|x)}].$ (32)
This loss is very simple and intuitive, term $-I\pi^{p}_{\theta}(\mathfrak{a}|x)$ try to maximize preferences, $\beta$ controls the strength of the regularization on SFT dataset, and $\gamma$ controls the strength of the regularization pretraining dataset.
Eliminate both the need for reward model and $\pi_{\text{ref}}$
By using preference as reward, we don’t need a reward model to approximate preference probability. By replacing KL-regularization $\mathbb{D}_{\text{KL}}[\pi^{s}_{\theta}||\pi^{s}_{\text{ref}}]$ with offline dataset regularization $-\mathbb{E}_{\langle x,y\rangle\sim\mathcal{D}_{\text{ref}}}[\log\pi^{s}_{\theta}(y|x)]$, we don’t need the reference policy $\pi^{s}_{\text{ref}}$. So, MPO algorithm simplify the learning process and saves memory usage.